OpenStudy (anonymous):

what is k divided by 24 = 85

OpenStudy (anonymous):

k=85*24?

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

multiply 85 times 24 and your answer to that equals k

OpenStudy (anonymous):

thnks whts 42=w/18

OpenStudy (anonymous):

42*18 = 756 so w = 756

OpenStudy (anonymous):

so antything lik tht problem u divide?

OpenStudy (anonymous):

so 255=x over 4 x overr 4 x 4= 255 x 255 x= 1020?

OpenStudy (anonymous):



OpenStudy (anonymous):

4(255)=x

OpenStudy (anonymous):

you do the opposite of what the equation is doing so in the first problem k/24 = 85 it is dividing so you multiply. in say x* 18 = 200 it is multipying so you divide

OpenStudy (anonymous):

don't foget to check your answer always just substitute your answer in the problem

OpenStudy (anonymous):

yep, thnk my teach always told us tht

OpenStudy (anonymous):

whats the answer tho